RIZZIO TECH IS A FOR- PROFIT VOCATIONAL SCHOOL. THE SCHOOL BASES ITS BUDGETS ON TWO MEASURES OF ACTIVITY (I.E. COST DRIVERS), NAMELY STUDENT AND COURSE. THE SCHOOL USES THE FOLLOWING DATA IN ITS BUDGETING:

FIXED ELEMENT PER MONTH VARIABLE ELEMENT PER STUDENT VARIABLE ELEMENT PER COURSE

REVENUE $0 $416 $0

FACULTY WAGES $0 $0 $2,300

COURSE SUPPLIES $0 $77 $25.00

ADMINISTRATIVE EXPENSES $57,900 $16 $27

IN SEPTEMBER, THE SCHOOL BUDGETED FOR 1,820 STUDENTS AND 173 COURSES. THE ACTUAL ACTIVITY FOR THE MONTH WAS 2,020 STUDENTS AND 178 COURSES.

REQUIRED:

PREPARE THE SCHOOL'S PLANNING BUDGET FOR SEPTEMBER
Answer #1

The planning budget will be prepared as follows:

Planning Budget
Revenue 757120
Expenses:
     Faculty wages 397900
     Course supplies 144465
     Administrative expenses 91691
Total expenses 634056
Net operating income 123064

Above figures have been calculated as follows:

Planning Budget
Revenue =416*1820
Expenses:
     Faculty wages =2300*173
     Course supplies =(77*1820)+(25*173)
     Administrative expenses =57900+(16*1820)+(27*173)
Total expenses =397900+144465+91691
Net operating income =757120-634056